# Solute Transport due to Periodic Loading in a Soft Porous Material

### v1

@@ Abstract
In soft porous media, deformation drives solute transport via the intrinsic coupling between flow of the fluid and rearrangement of the pore structure. Solute transport driven by periodic loading, in particular, can be of great relevance in applications ranging from the geomechanics of contaminants in the subsurface to the biomechanics of nutrient transport in living tissues, scaffolds for tissue engineering, and biomedically employed hydrogels. However, the basic features of this process have not previously been systematically investigated. Here, we fill this hole in the context of a 1D model problem. We do so by expanding the results from a companion study, in which we explored the poromechanics of periodic deformations, by introducing and analysing the impact of the resulting fluid and solid motion on solute transport. We first characterise the independent roles of the three main mechanisms of solute transport in porous media — advection, molecular diffusion, and hydrodynamic dispersion — by examining their impacts on the solute concentration profile during one loading cycle. We next explore the impact of the transport parameters, showing how these alter the relative importance of diffusion and dispersion. We then explore the loading parameters by considering a range of loading periods — from slow to fast, relative to the poroelastic timescale — and amplitudes — from infinitesimal to large. We show that solute spreading over several loading cycle increases monotonically with amplitude, but is maximised for intermediate periods because of the increasing poromechanical localisation of the flow and deformation near the permeable boundary as the period decreases.
@@ I Introduction
Solute transport in deformable porous media has been studied previously in several different contexts. The impact of soil consolidation on transport has been studied primarily for geomechanical applications (e.g., Fox, 2007a, b; Pu et al., 2018). In that context, it has been shown that solute transport is enhanced by larger deformations Peters and Smith (2002); Lewis et al. (2009); Alshawabkeh and Rahbar (2006); Zhang et al. (2013); Bonazzi et al. (2021). Deformation has also been shown to increase mixing and reduce breakthrough time in the context of miscible viscous fingering Tran and Jha (2020). The key feature introduced by periodic loading is the continuously fluctuating fluid flow, which can irreversibly modify diffusion and dispersion even when the macroscopic advective component is perfectly reversible. The role of periodic flow in enhancing solute transport and mixing has been studied in rigid and compressible 1D pore networks Goldsztein and Santamarina (2004); Claria et al. (2012). In a poroelastic material, the impact of small periodic deformations on solute transport has been analysed for a range of frequencies Bonazzi et al. (2021). In heterogeneous coastal aquifers, periodic flow due to tidal loading has been shown to enhance solute transport Wu et al. (2020); Trefry et al. (2019). However, all of these studies examined a specific and relatively limited range of parameters.
@@
In the context of biological tissues, periodic loading is known to enhance the transport of nutrients (Ferguson et al., 2004; Gardiner et al., 2007; Witt et al., 2014; Schmidt et al., 2010; Zhang, 2011; DiDomenico et al., 2017; Zhang and Szeri, 2008). Similarly, periodic deformations are used to enhance the infiltration of solutes into hydrogels Albro et al. (2008); Vaughan et al. (2013) and other scaffolds for tissue engineering Mauck et al. (2003); Cortez et al. (2016); Kumar et al. (2018); Fan et al. (2016), where the correlation between loading parameters, nutrient transport and cell survival is of particular interest. Increasing the loading amplitude and/or decreasing the loading period induces a transition from diffusion-dominated to advection-dominated regimes Urciuolo et al. (2008) and amplifies the role of hydrodynamic dispersion Sengers et al. (2004). Decreasing the loading period also leads to localisation of flow and deformation near permeable boundaries, resulting in larger velocities near the surface that promote external solute infiltration Gardiner et al. (2007); Urciuolo et al. (2008); DiDomenico et al. (2017); Vaughan et al. (2013).
@@
In general, despite the established role of hydrodynamic dispersion in driving the transport of solutes in porous media, dispersion is rarely included in biomechanical models (with the notable exception of Ref. Sengers et al. (2004)). One context where dispersion is widely agreed to be important is in brain microcirculation Kelley and Thomas (2023). In the vascular network within the brain, dispersion results from the shear-induced radial concentration gradients in single vessels (e.g. Marbach and Alim (2019); Sharp et al. (2019); Berg et al. (2020); Troyetsky et al. (2021); Bojarskaite et al. (2023)) and the progressive bifurcation of vessels into smaller branches that can be modelled at the continuum scale as a porous material (e.g. Zimmerman and Tartakovsky (2020); Goirand et al. (2021)).
@@
Dispersion is typically neglected in the context of tissues and gels for two main reasons. First, fluid flow is often assumed to be slow, implying that transport is dominated by diffusion. In other words, the Peclet number $\mathrm{Pe}=VL/D$ is assumed to be small, where $V$ is the fluid velocity, $L$ the streamwise length scale, and $D$ the diffusivity. However, it is straightforward to show that $\mathrm{Pe}$ can be of order 1 or larger in a tissue or gel subject to fast ($0.1-1$ Hz) and large ($10-20\%$) deformations (see table II), suggesting that dispersion may be important or even dominant in some scenarios Delgado (2007). Indeed, many studies highlight a transition from diffusion-dominated to advection-dominated transport without acknowledging the potential role of dispersion Gardiner et al. (2007); Urciuolo et al. (2008); DiDomenico et al. (2017); Vaughan et al. (2013). With an analogous argument, Davit et al. (2013) illustrated the importance of including dispersion in models for solute transport in biofilms. The second typical reason for neglecting dispersion in tissues and gels is the assumption that the longitudinal and transverse dispersivities themselves are negligible. This expectation is a result of physical insight derived from transport in granular materials, where the dispersivity is typically taken to be proportional to the pore size Saffman (1959); Oswald and Kinzelbach (2004); Kree and Villermaux (2017); Liang et al. (2018). Indeed, the typical pore size is $\sim 10$ nm in polymeric gels and in the extra-cellular matrix of tissues (e.g., around 6 nm in cartilage Mow et al. (1984)) and can therefore be similar to (or smaller than) the size of large solute molecules Maroudas (1970); DiDomenico et al. (2018), originating solid-solute friction Ateshian and Weiss (2013); Yao and Gu (2007). However, tissues and scaffolds are heterogeneous and multiscale materials; the presence of other components, such as collagen fibres, originates a “mesoscale” of larger pores (e.g., 100–150 nm in cartilage Maroudas (1975); Levick (1987); Federico and Herzog (2008)), where even larger solute molecules can pass DiDomenico et al. (2017, 2018) and where dispersion is likely to play a much larger role. The same is true for double-porosity scaffolds and gels, where additional channels and/or macroscopic pores are included to enhance fluid flow throughout the scaffold depth Buijs et al. (2010); Lee et al. (2015); Mesallati et al. (2013). As a further counter-argument, we hypothesise that, even in pores that are small compared to the solute molecules, the irrelevance of pore-scale velocity gradients does not exclude velocity variations and streamline alterations in the overall network, which could originate longitudinal and transverse dispersion. This hypothesis is consistent with the quantification of tortuosity in several soft tissues Maroudas (1970); Hrabe et al. (2004); Zhang and Szeri (2005).
@@
Thus, the impact of periodic loading on solute transport in soft porous media has been addressed with various approaches and assumptions across a variety of specific applications in soils, tissues, hydrogels, and scaffolds. However, no single study has yet provided a comprehensive understanding across a wide range of loading frequencies and amplitudes. Moreover, the impact of hydrodynamic dispersion remains relatively unexplored and therefore poorly understood, particularly in the context of biological and biomedical applications. Here, we study the transport and mixing of solutes due to arbitrarily large, periodic deformations of a soft porous material. For the flow and deformation, we adopt a one-dimensional, large-deformation poroelasticity model that includes rigorous nonlinear kinematics, deformation-dependent permeability, and Hencky elasticity for the solid skeleton. We used this model to explore the poromechanics of large-amplitude periodic loading in a companion study (Fiori et al., 2023). Here, we introduce a model for solute transport due to advection, diffusion, and dispersion. We study the separate roles of advection, diffusion, and dispersion during one loading cycle. We then consider the impact of the transport and loading parameters on transport and mixing over longer time periods and/or larger number of loading cycles. We report the impact of a wide range of loading amplitudes and periods on each transport mechanism and observe how transport depends on the poromechanical response through its impact on local fluid flow. When dispersion is negligible, we show that diffusion is insensitive to loading period but slightly suppressed by increased loading amplitude. With dispersion, larger amplitudes always boost solute spreading; however, progressively shorter periods impact transport and mixing in more complex ways: fast loading promotes spreading by inducing large fluid velocities, but very fast loading hinders spreading by progressively localising the flow and deformation. We show that the competition between these two effects results in maximum solute transport and mixing for intermediate loading periods.
@@ II.2 Kinematics
We consider an Eulerian reference frame, in which the solid displacement is $\mathbf{u_{s}}=\mathbf{x}-\mathbf{X}(\mathbf{x},t)$, with $\mathbf{X}(\mathbf{x},t)$ the reference position of the material point that at time $t$ occupies position $\mathbf{x}$. We choose our reference configuration to be the relaxed configuration, such that $\mathbf{X}(\mathbf{x},0)=\mathbf{x}$ and $\mathbf{u_{s}}(\mathbf{x},0)=0$. The true volume fractions of fluid and solid are $\phi_{f}$ and $\phi_{s}$, respectively, where $\phi_{f}+\phi_{s}=1$. In this uniaxial setting, the solid displacement and the solid and fluid velocities are one-dimensional and given by
@@ II.4 Mechanical equilibrium and elasticity law
where $\sigma^{\prime}$ is the $xx$ component of $\bm{\sigma}$.
@@
is the poroelastic diffusivity.
We use Hencky hyperelasticity Hencky (1931) to capture kinematic nonlinearity, as discussed in detail in Ref. Fiori et al. (2023). For a uniaxial deformation, the relevant component of the effective stress is then MacMinn et al. (2016); Auton and MacMinn (2018)
@@
where $\mathcal{M}$ is the $p$-wave or oedometric modulus. With appropriate initial and boundary conditions and the normalised Kozeny-Carman permeability law, equations (11), (12), and (13) comprise a closed model for the evolution of the porosity.
@@ II.5 Solute transport
In 1D, conservation of mass for the solute can be written
@@
The first term in the square brackets is the Darcy-scale solute flux due to advection, which occurs here entirely in response to the deformation. The second term in the square brackets combines molecular diffusion and hydrodynamic dispersion, thus taking the latter to be a Fickian process (e.g. Scheidegger (1961)). The latter term is multiplied by the porosity $\phi_{f}$ since solute movements only occur in the fluid phase. The coefficient $\mathcal{D}$ can be written
@@
where $\mathcal{D}_{m}$ and $\mathcal{D}_{h}$ are the coefficients of molecular diffusion and hydrodynamic dispersion, respectively. Dispersion, in which pore-scale velocity gradients and the tortuosity of the pore space lead to macroscopic spreading of solute, depends sensitively on flow conditions and the details of the pore structure in ways that are not yet fully understood, even for rigid porous materials Dentz et al. (2018, 2023). The most widely used model for the macroscopic dispersive flux is Fickian, as above, with a velocity-dependent dispersion coefficient given in 1D by Scheidegger (1961); Brenner and Edwards (1993); Whitaker (1998); Gelhar (1993)
@@
where $\alpha$ is the longitudinal dispersivity. Note that the dispersive flux is therefore proportional to $|v_{f}-v_{s}|$, unlike the advective flux, because dispersion is driven by flow of fluid through the pore structure (i.e., $v_{f}=v_{s}\neq 0$ would lead to advection but no dispersion). Note also that unlike the advective flux, the diffusive and the dispersive fluxes are independent of the direction of the fluid flow. The dispersivity $\alpha$ is typically taken to be a constant material property for a given pore structure. In a deforming porous material, and particularly for moderate to large deformations, it is likely that $\alpha$ should be deformation-dependent to account for the evolving pore structure. This coupling has not yet been studied and will be the subject of future work. Here, we take $\alpha$ to be constant for simplicity.

### v2

@@ Abstract
In soft porous media, deformation drives solute transport via the intrinsic coupling between flow of the fluid and rearrangement of the pore structure. Solute transport driven by periodic loading, in particular, can be of great relevance in applications including the geomechanics of contaminants in the subsurface and the biomechanics of nutrient transport in living tissues and scaffolds for tissue engineering. However, the basic features of this process have not previously been systematically investigated. Here, we fill this hole in the context of a 1D model problem. We do so by expanding the results from a companion study, in which we explored the poromechanics of periodic deformations, by introducing and analysing the impact of the resulting fluid and solid motion on solute transport. We first characterise the independent roles of the three main mechanisms of solute transport in porous media — advection, molecular diffusion, and hydrodynamic dispersion — by examining their impacts on the solute concentration profile during one loading cycle. We next explore the impact of the transport parameters, showing how these alter the relative importance of diffusion and dispersion. We then explore the loading parameters by considering a range of loading periods — from slow to fast, relative to the poroelastic timescale — and amplitudes — from infinitesimal to large. We show that solute spreading over several loading cycle increases monotonically with amplitude, but is maximised for intermediate periods because of the increasing poromechanical localisation of the flow and deformation near the permeable boundary as the period decreases.
@@ I Introduction
Solute transport in deformable porous media has been studied in several different contexts. The impact of monotonic soil consolidation on transport has been studied extensively for its relevance to the management of landfills and other contaminated sediments, such as dredging or mining waste (e.g., Smith, 2000; Peters and Smith, 2002; Alshawabkeh and Rahbar, 2006; Fox, 2007a, b; Arega and Hayter, 2008; Lewis et al., 2009; Zhang et al., 2012, 2013; Xie et al., 2016; Pu et al., 2018; Bonazzi et al., 2021). In that context, it is well known that consolidation enhances solute transport. Deformation has also been shown to increase mixing and reduce breakthrough time in the context of miscible viscous fingering Tran and Jha (2020). The key feature introduced by periodic loading is the continuously fluctuating fluid flow, which can irreversibly modify diffusion and dispersion even when the macroscopic advective component is perfectly reversible. The role of periodic flow in enhancing solute transport and mixing has been studied in rigid and compressible 1D pore networks Goldsztein and Santamarina (2004); Claria et al. (2012). In a poroelastic material, solute transport due to small periodic deformations has been explored across a range of parameters, including compressibility and forcing frequency, for semi-infinite homogeneous systems Pool et al. (2016), finite homogeneous systems Bonazzi et al. (2021), and finite heterogeneous systems Trefry et al. (2019); Wu et al. (2020). The latter two studies focus in particular on the combined role of poroelasticity, heterogeneity, and transient forcing in generating chaotic advection.
@@
Periodic loading is also known to enhance the transport of nutrients in biological tissues (Ferguson et al., 2004; Gardiner et al., 2007; Witt et al., 2014; Schmidt et al., 2010; Zhang, 2011; DiDomenico et al., 2017b; Zhang and Szeri, 2008). Similarly, periodic deformations are used to enhance the infiltration of solutes into hydrogels Albro et al. (2008); Vaughan et al. (2013) and other scaffolds for tissue engineering Mauck et al. (2003); Cortez et al. (2016); Kumar et al. (2018); Fan et al. (2016), where the correlation between loading parameters, nutrient transport, and cell survival is of particular interest. Increasing the loading amplitude and/or decreasing the loading period induces a transition from diffusion-dominated to advection-dominated regimes Urciuolo et al. (2008) and amplifies the role of hydrodynamic dispersion Sengers et al. (2004). Decreasing the loading period also leads to localisation of flow and deformation near permeable boundaries, resulting in larger velocities near the surface that promote external solute infiltration Gardiner et al. (2007); Urciuolo et al. (2008); DiDomenico et al. (2017b); Vaughan et al. (2013).
@@
In general, despite the established role of hydrodynamic dispersion in driving the transport of solutes in porous media, dispersion is rarely included in biomechanical models (with the notable exception of Ref. Sengers et al. (2004)). One context where dispersion is widely agreed to be important is in brain microcirculation Kelley and Thomas (2023). In the vascular network within the brain, dispersion results from the shear-induced radial concentration gradients in single vessels (e.g., Marbach and Alim, 2019; Sharp et al., 2019; Berg et al., 2020; Troyetsky et al., 2021; Bojarskaite et al., 2023) and the progressive bifurcation of vessels into smaller branches that can be modelled at the continuum scale as a porous material (e.g., Zimmerman and Tartakovsky, 2020; Goirand et al., 2021).
@@
Dispersion is typically neglected in the context of tissues and gels for two main reasons. First, fluid flow is often assumed to be slow, implying that transport is dominated by diffusion. In other words, the Péclet number $\mathrm{Pe}=VL/\mathcal{D}_{m}$ is assumed to be small, where $V$ is the characteristic fluid velocity, $L$ the characteristic streamwise length scale, and $\mathcal{D}_{m}$ the molecular diffusivity. However, it is straightforward to show that $\mathrm{Pe}$ can be order 1 or larger in a tissue or gel subject to fast (0.10.10.10.11111 Hz) and large (1010101020202020%) deformations (see table II), suggesting that dispersion may be important or even dominant in some scenarios Delgado (2007). Indeed, many studies highlight a transition from diffusion-dominated to advection-dominated transport without acknowledging the potential role of dispersion Gardiner et al. (2007); Urciuolo et al. (2008); DiDomenico et al. (2017b); Vaughan et al. (2013). With an analogous argument, Davit et al. (2013) illustrated the importance of including dispersion in models for solute transport in biofilms. The second typical reason for neglecting dispersion in tissues and gels is the assumption that the longitudinal and transverse dispersivities themselves are negligible. This expectation is a result of physical insight derived from transport in granular materials, where the dispersivity is typically taken to be proportional to the pore size Saffman (1959); Oswald and Kinzelbach (2004); Kree and Villermaux (2017); Liang et al. (2018). Indeed, the typical pore size is $\sim 10$ nm in polymeric gels and in the extra-cellular matrix of tissues (e.g., around 6 nm in cartilage Mow et al. (1984)) and can therefore be similar to (or smaller than) the size of large solute molecules Maroudas (1970); DiDomenico et al. (2018), originating solid-solute friction Ateshian et al. (2011); Yao and Gu (2007). However, tissues and scaffolds are heterogeneous and multiscale materials; the presence of other components, such as collagen fibres, originates a “mesoscale” of larger pores (e.g., 100–150 nm in cartilage Maroudas (1975); Levick (1987); Federico and Herzog (2008)), where even larger solute molecules can pass DiDomenico et al. (2017a, 2018) and where dispersion is likely to play a much larger role. The same is true for double-porosity scaffolds and gels, where additional channels and/or macroscopic pores are included to enhance fluid flow throughout the scaffold depth Op Den Buijs et al. (2010); Lee et al. (2015); Mesallati et al. (2013). As a further counter-argument, we hypothesise that, even in pores that are small compared to the solute molecules, the irrelevance of pore-scale velocity gradients does not exclude velocity variations and streamline alterations in the overall network, which could originate longitudinal and transverse dispersion. This hypothesis is consistent with the quantification of tortuosity in several soft tissues Maroudas (1970); Hrabe et al. (2004); Zhang and Szeri (2005).
@@
Thus, the impact of periodic loading on solute transport in soft porous media has been addressed with various approaches and assumptions across a variety of specific applications in soils, tissues, hydrogels, and scaffolds. However, no single study has yet provided a comprehensive understanding across a wide range of loading frequencies and amplitudes. Moreover, the impact of hydrodynamic dispersion remains relatively unexplored and therefore poorly understood, particularly in the context of biological and biomedical applications. Here, we study the transport and mixing of solutes due to arbitrarily large, periodic deformations of a soft porous material. For the flow and deformation, we adopt a one-dimensional, large-deformation poroelasticity model that includes rigorous nonlinear kinematics, deformation-dependent permeability, and Hencky elasticity for the solid skeleton. In a companion study, we used this model to explore the poromechanics of large-amplitude periodic loading (Fiori et al., 2023). Here, we additionally consider solute transport due to advection, diffusion, and dispersion. We first study the separate roles of advection, diffusion, and dispersion during one loading cycle. We then consider the impact of the transport and loading parameters on transport and mixing over longer time periods and/or larger numbers of loading cycles. We report the impact of a wide range of loading amplitudes and periods on each transport mechanism and observe how transport depends on the poromechanical response through its impact on local fluid flow. When dispersion is negligible, we show that diffusion is insensitive to loading period but slightly suppressed by increased loading amplitude. With dispersion, larger amplitudes always boost solute spreading; however, progressively shorter periods impact transport and mixing in more complex ways: fast loading promotes spreading by inducing large fluid velocities, but very fast loading hinders spreading by progressively localising the flow and deformation. We show that the competition between these two effects results in maximum solute transport and mixing for intermediate loading periods.
@@ II.2 Kinematics
We consider an Eulerian reference frame, in which the solid displacement is $\mathbf{u_{s}}=\mathbf{x}-\mathbf{X}(\mathbf{x},t)$, with $\mathbf{X}(\mathbf{x},t)$ the reference position of the solid material point that at time $t$ occupies position $\mathbf{x}$. We choose our reference configuration to be the relaxed configuration, such that $\mathbf{X}(\mathbf{x},0)=\mathbf{x}$ and $\mathbf{u_{s}}(\mathbf{x},0)=0$. The true volume fractions of fluid and solid are $\phi_{f}$ and $\phi_{s}$, respectively, where $\phi_{f}+\phi_{s}=1$. In this uniaxial setting, the solid displacement and the solid and fluid velocities are one-dimensional and given by
@@ II.4 Mechanical equilibrium and elasticity law
where $\sigma^{\prime}$ is the $xx$ component of $\bm{\sigma}^{\prime}$.
@@
is the poroelastic diffusivity. Note that a very similar model is used for the solidification of colloidal suspensions in applications such as filtration and sedimentation, for which the poroelastic diffusivity $D_{f}(\phi_{f})$ (i.e., the “solids diffusivity”) is characterised as a composite material property (e.g., Davis and Russel, 1989; Peppin et al., 2006; Style and Peppin, 2011; Bouchaudy and Salmon, 2019; Worster et al., 2021).
We use Hencky hyperelasticity Hencky (1931) as a simple, large-deformation model that captures kinematic nonlinearity. For a uniaxial deformation, the relevant component of the effective stress is then
@@
where $\mathcal{M}$ is the $p$-wave or oedometric modulus (MacMinn et al., 2016). Note that, for these constitutive choices of Kozeny-Carman permeability and Hencky elasticity, the permeability at the left boundary can vanish for sufficiently large $A$ and/or small $T$, because the poroelastic diffusivity remains finite rather than diverging as $\phi_{f}\to 0$ (Hewitt et al., 2016). We motivate and discuss our constitutive choices and explore the poroelastic diffusivity in more detail in Ref. Fiori et al. (2023).
With appropriate initial conditions, boundary conditions, and the normalised Kozeny-Carman permeability law, equations (11), (12), and (13) comprise a closed model for the evolution of the porosity.
@@ II.5 Solute transport
In 1D, it is well known that conservation of mass for a passive solute can be written
@@
The first term in the square brackets is the Darcy-scale solute flux due to advection, which occurs here entirely in response to the deformation. The second term in the square brackets combines molecular diffusion and hydrodynamic dispersion, thus taking the latter to be a Fickian process (e.g., Scheidegger, 1961). The latter term is multiplied by the porosity $\phi_{f}$ since solute movements only occur in the fluid phase. The coefficient $\mathcal{D}$ can be written
@@
where $\mathcal{D}_{m}$ and $\mathcal{D}_{h}$ are the coefficients of molecular diffusion and hydrodynamic dispersion, respectively. Dispersion, in which pore-scale velocity gradients and the tortuosity of the pore space lead to macroscopic spreading of solute, depends sensitively on flow conditions and the details of the pore structure in ways that are not yet fully understood, even for rigid porous materials (Dentz et al., 2018, 2023). The most widely used model for the macroscopic dispersive flux is Fickian, as above, with a velocity-dependent dispersion coefficient given in 1D by
@@
where $\alpha$ is the longitudinal dispersivity (Scheidegger, 1961; Brenner and Edwards, 1993; Whitaker, 1999; Gelhar, 1993). Note that the dispersive flux is proportional to $|v_{f}-v_{s}|$, unlike the advective flux, because dispersion is driven by flow of fluid through the pore structure (i.e., $v_{f}=v_{s}\neq 0$ would lead to advection but no dispersion). Note also that, unlike the advective flux, the diffusive and dispersive fluxes are independent of the direction of the fluid flow.
The dispersivity $\alpha$ is typically taken to be a constant material property for a given pore structure. In a deforming porous material, and particularly for moderate to large deformations, it is likely that $\alpha$ should be deformation-dependent to account for the evolving pore structure. For example, particle-particle interactions and rearrangements are known to drive enhanced dispersion in dense suspensions Souzy et al. (2016, 2017) and compaction has been shown to have a nontrivial impact on dispersion in bead packs and packed beds Charlaix et al. (1987); Östergren and Trägårdh (2000); Liu et al. (2024). We expect similar but even larger effects in poroelastic materials under large deformations, which may ultimately require novel dispersion models, but these phenomena are beyond the scope of the present study. Here, we take $\alpha$ to be a constant for simplicity.